a diagonal of a rectangle is 60 cm long .if the length of the rectangle is 12 cm greater than the width .find the dimensions

Answer by mananth(16946)

width -------x
length ---------x+12
diagonal --------60
They form a right triangle
x^2+(x+12)^2=60^2
x^2+x^2+24x+144=3600
2x^2+24x-3456=0
/2
x^2+12x-1728=0
x^2+48x-36x-1728=0
x(x+48)-36(x+48)=0
(x+48)(x-36)=0
x=36 which is positive
36 & 48 are the sides